Find the x-intercept(s) and the coordinates of the vertex for the parabola y = x^2-2x-8. If there is more than one x-intercept, separate them with commas.

Answer by lwsshak3(11628)

Find the x-intercept(s) and the coordinates of the vertex for the parabola y = x^2-2x-8. If there is more than one x-intercept, separate them with commas.
**
y = x^2-2x-8
complete the square
y = (x^2-2x+1)-8-1
y = (x-1)^2-9
vertex: (1,-9)
x-intercepts
set y=0
(x-1)^2-9=0
(x-1)^2=9
(x-1)=±√=±√9=±3
x=-2,4